OpenStudy (anonymous):

Solve. 2x2 – 6x = –8

OpenStudy (anonymous):

by completing the square

OpenStudy (anonymous):

2x² - 6x = -8 2x² - 6x + 8 = 0 /2 x² - 3x + 4 = 0 \[x = \frac{ -3 \pm \sqrt{9 - 16} }{ 2 }\] \[x = \frac{ -3 \pm \sqrt{-5} }{ 2 }\] At this point you got a negative square root and to solve it you have to use the complex numbers that I don't remember how to use :/